### v1

# Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems

Abstract
A quantum kinetic theory of the linear response to an electric field is provided from a controlled expansion of the Keldysh theory at leading order, for a general multiband electron system with weak scalar disorder. The response is uniquely partitioned into intraband distribution functions and interband quantum coherences. A new explicit formula is provided for the latter, outlining their true nature as local slave quantities, for which the mesoscopic gradients of the former is uncovered as a new source term. The precise connection with the Kubo formula with vertex corrections in the ladder approximation is established. A pedagogical application to spin-orbit torque theory in the two dimensional electron gas demonstrates the striking efficiency of the new formalism. Some important implications for the theory of orbital transport are also briefly discussed.
Introduction– Linear response theory based on the Kubo formula[1, 2], usually expressed in more practical forms of the Kubo-Bastin or Kubo-Streda types [3, 4, 5, 6, 7], is instrumental to our understanding of the electron Fermi liquid properties in solids. Its application to the theory of quantum Hall effect [8, 9] provides a salient illustration of its power in this context. Berry phase effects [10] are fully included, as seen for instance in the expressions one can derive for the intrinsic anomalous Hall effect (AHE) [11], spin Hall effect (SHE) [12] or orbital Hall effect (0HE) [13, 14] conductivities. It is also naturally amenable to diagrammatic expansions [15, 16, 17], allowing to account for disorder in a controlled manner. However, concrete Kubo based calculations often incorporate disorder effects only through a constant broadening approximation (CBA) [18, 19]. Such an uncontrolled approximation, completely neglecting vertex corrections, is often problematic [20, 21]. The Kubo formula also enables direct computations from first principles, see for instance [22, 23, 24, 25], leading to quantitative predictions for actual material systems. However, in many systems under current intense investigation, such as spintronic [26], orbitronic [27] or topological [28] nanostructured materials and devices, one is primarily interested in disorder averaged responses in finite mesoscopic samples. Full quantum simulations in such cases, specially if one wants to go beyond the CBA, become quite involved, even with the emergence of efficient real space numerical formulations [29, 30, 31, 32, 33]. Alternatively, linear response can be derived from semiclassical approaches [34, 35, 36, 37], which can be seen as some attempts to incorporate leading quantum geometric effects and disorder corrections to the simple quasiparticle transport picture provided by the Boltzmann equation (BE) [38]. While this is in principle more suited to model spatially inhomogeneous systems, these earlier works did not provide a full account of the effects of gradients on mesoscopic length scales, while incorporating on an equal footing leading Berry phase effects and vertex corrections associated with weak disorder.
In this letter, we provide a consistent quantum kinetic theory of the linear response to a nearly constant and uniform electric field ${\bm{E}}$, for a multiband electron system modeled by a low energy Hamiltonian, in presence of weak scalar disorder. This has been obtained from a controlled semiclassical expansion of the Schwinger-Keldysh theory [39, 40, 41], in a gauge invariant formulation in phase space [42] now extended beyond the clean limit. With the details of all the derivations to be found here [43], we first expose our general results in terms of intraband and interband responses, uncovering mesoscopic gradients of the former as a previously overlooked source of the latter. With the help of the Matsubara formalism [15, 16], we then establish the equivalence of our results with the Kubo formula with vertex correction in the ladder approximation, for the spatially homogeneous case. As a pedagogical example, we consider anew the theory of spin-orbit torque (SOT) in the two-dimensional electron gas (2DEG), and uncover a new instance of intrinsic response as a Fermi surface property, reminiscent of the classical result obtained by Haldane in the case of the AHE [44]. In conclusion, we briefly discuss future extensions and we outline some implications for the theory of the recently observed orbital edge accumulation [45, 46] and orbital torque [47, & references therein].
General formalism – Our main new result is a general expression for the interband part of the density matrix i.e., $\hat{\rho}_{inter}^{(1)}=\sum_{n\neq m}\hat{\rho}_{nm}^{(1)}$, in linear response to ${\bm{E}}$, with
$\hat{\rho}_{nm}^{(1)}$ $=-\hbar\frac{f_{n}^{(0)}-f_{m}^{(0)}}{\varepsilon_{n}-\varepsilon_{m}}e{\bm{E}}\cdot\hat{\mathcal{\bm{A}}}_{nm}-\hbar\partial_{\bm{x}}\frac{f_{n}^{(1)}+f_{m}^{(1)}}{2}\cdot\hat{\mathcal{\bm{A}}}_{nm}$
$\!\!\!\!\!\!\!-\frac{{\bf i}\hbar}{2(\varepsilon_{n}-\varepsilon_{m})}\frac{2\pi n_{i}v_{0}^{2}}{\hbar}\sum_{q}\int\!\!\frac{d^{d}\underaccent{\tilde}{p}}{(2\pi\hbar)^{d}}\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\hat{P}_{m}\times\cdots$ (1)
×$\!\!\!\!\!\!\!\!\!\!\!\cdots\times\left[\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{n}\right)\left(\underaccent{\tilde}{f}_{q}^{(1)}\!\!-\!\!f_{n}^{(1)}\right)+\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{m}\right)\left(\underaccent{\tilde}{f}_{q}^{(1)}\!\!-\!\!f_{m}^{(1)}\right)\right],$
while the intraband part $\hat{\rho}_{intra}^{(1)}=\sum_{n}f^{(1)}_{n}\hat{P}_{n}$ is controlled by a system of linear BEs
$\partial_{t}f_{n}^{(1)}+\partial_{\bm{p}}\varepsilon_{n}\cdot\partial_{\bm{x}}f_{n}^{(1)}+(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{n}\partial_{\varepsilon}f_{n}^{(0)}=C_{n}[{\bm{f}}^{(1)}].$ (2)
To be more precise, $\hat{\rho}^{(1)}({\bm{x}},t;{\bm{p}})\equiv\hat{\rho}^{(1)}=\hat{\rho}^{(1)}_{intra}+\hat{\rho}^{(1)}_{inter}$ as determined by Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) is the gauge invariant Wigner function i.e., the gauge invariant Wigner transform of the density matrix [48]. Once $\hat{\rho}^{(1)}$ is obtained, the expectation value of any single particle observable $\smash{{\hat{O}}}$, in the form of its local density $\smash{o^{(1)}({\bm{x}},t)}$, is obtained from its Weyl symbol $\smash{{\hat{o}}({\bm{x}},{\bm{p}})}$ as $o^{(1)}({\bm{x}},t)=\int\frac{d^{d}p}{(2\pi\hbar)^{d}}\ \mathrm{Tr}[{\hat{\rho}}^{(1)}({\bm{x}},t;{\bm{p}})\ {\hat{o}}({\bm{x}},{\bm{p}})]$. Equations (Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) determine the linear response of a non interacting system of electrons with charge $e(<0)$, specified by a multiband low energy Hamiltonian matrix i.e., ${\hat{h}}={\hat{h}}({\bm{p}})$, with ${\bm{p}}$ the kinetic crystal pseudo momentum. This matrix is acting on the Hilbert space of dimension $N$ of the internal degrees of freedom (spin, orbital), and it may equally correspond to a “toy” model or to a sophisticated Wannier tight-binding Hamiltonian [49, 50, 51]. It has in general $M\leq N$ distinct eigenvalues, the band energies, introduced as $\varepsilon_{n}\equiv\varepsilon_{n}({\bm{p}})$. The projectors into the corresponding eigenspaces, introduced as $\hat{P}_{n}\equiv\hat{P}_{n}({\bm{p}})$, read [52] $\hat{P}_{n}=\prod_{m}\frac{\hat{h}-\varepsilon_{m}\hat{\gamma}_{0}}{\varepsilon_{n}-\varepsilon_{m}}$, with $\hat{\gamma}_{0}$ the identity matrix. We further define $\hat{\mathcal{\bm{A}}}_{nm}={\bf i}\hat{P}_{n}\partial_{\bm{p}}\hat{P}_{m}$, which appears in the first line of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). In the special case of two non degenerate bands, it is immediate to verify that we have $(\hat{\mathcal{\bm{A}}}_{nm})^{i}={\mathcal{A}}^{i}_{nm}|n\rangle\langle m|$, with ${\mathcal{A}}^{i}_{nm}={\bf i}\langle n|\partial_{p^{i}}m\rangle$ the previously introduced abelian interband Berry connection [53]. Hence, we recognize $\hat{\mathcal{\bm{A}}}_{nm}$ as a generalized non abelian interband Berry connection, defined in a manifestly $SU(N)$ gauge invariant way. Also appearing in Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2), are the linear changes in the intraband distribution functions i.e., $f_{n}^{(1)}\equiv f_{n}^{(1)}({\bm{x}},t;{\bm{p}})$, reducing in equilibrium to the Fermi-Dirac distribution function $f_{n}^{(0)}\equiv n_{\scriptscriptstyle FD}(\varepsilon_{n})$. In the present work, we purposely limit our consideration of disorder to the most simple and generic model i.e., a dilute random distribution of scalar point scattering potentials treated at the level of the self-consistent second Born approximation after configuration averaging. It is parameterized by the density $n_{i}$ of defects and by a momentum independent scattering amplitude $v_{0}$. The collision integral appearing on the right-hand-side of the BE (2) reads
$\!\!\!\!C_{n}[{\bm{f}}^{(1)}]\!=\!\!\sum_{m}\!\!\int\!\!\!\frac{d^{d}\underaccent{\tilde}{p}}{(2\pi\hbar)^{d}}\delta\left(\underaccent{\tilde}{\varepsilon}_{m}-\varepsilon_{n}\right)W_{nm}\left(\underaccent{\tilde}{f}_{m}^{(1)}-f_{n}^{(1)}\right)\!,$ (3)
with the scattering kernel given by
$W_{nm}\equiv W_{nm}({\bm{p}},\underaccent{\tilde}{\bm p})=\frac{2\pi n_{i}v_{0}^{2}}{\hbar}{\rm Tr}\left[\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{m}\right].$ (4)
It has to be noted that we systematically use a tilde under-accent to indicate that a certain quantity depends on the integration momentum variable $\underaccent{\tilde}{\bm p}$. After introducing $\tau\sim\hbar/(n_{i}v_{0}^{2})$ as the momentum relaxation time scale, we see that $\xi_{i}=1/\tau$ defines a natural scaling parameter for the disorder, while $\hbar$ is of course the formal expansion parameter for the quantum corrections to the classical kinetics.
While the commonly considered decomposition of the Kubo-Streda type formulas into Fermi sea and Fermi surface terms [5, 7] is not physically significant in itself [44], the quantum kinetic approach leads naturally to a unique and gauge invariant split at the Wigner function level i.e., $\hat{\rho}^{(1)}=\hat{\rho}^{(1)}_{intra}+\hat{\rho}^{(1)}_{inter}$. Its physical meaning is further illuminated by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), establishing the interband response components as slave variables locally and uniquely determined by the quantum geometry and the solutions of the BE (2). Hence, our formalism establishes that direct interband contributions to observable densities are not quantities susceptible to advective or diffusive transport in real space. Further looking into the different terms on the right-hand side (RHS) of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we recognize the first term as of purely quantum geometric origin and of order $\hbar\xi_{i}^{0}$, which will yield for instance the intrinsic contribution to the AHE [54, 11] which can be expressed in terms of the Berry curvature. If we look now at the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), since $f_{n}^{(1)}\sim\xi_{i}^{-1}$, as it can be immediately verified by direct inspection of Eqs.(2-4), we see that it is also of order $\hbar\xi_{i}^{0}$ due to the compensating factor $(n_{i}v_{0}^{2})/\hbar\sim\xi_{i}$ in front of the momentum integral. Hence, even if this contribution is disorder mediated, it is ultimately independent of its strength. As demonstrated in the next section, this corresponds to the vertex correction in the ladder approximation to the interband part of the Kubo formula, for which we provide here a new explicit expression at the Wigner function level, in terms of the solutions to the BE (2). As previously mentioned, even recent theoretical studies rely most often on a simplified version of the Kubo-Streda formula at the CBA level, with the leading disorder correction to the interband response found to be of order $\hbar\xi_{i}$ [55], which is qualitatively wrong and totally misses the true leading effect of disorder. As for the second term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), it indicates that mesoscopic gradients in spatially inhomogeneous solutions of the BE (2) are also a source of interband quantum coherences, for which we provide here an explicit expression. To the best of our knowledge, this is an entirely new finding. Such gradients typically appear near boundaries and interfaces on the length scale of the electron mean-free-path [56, 57, 58], hence $\partial_{\bm{x}}f_{n}^{(1)}\sim\xi_{i}^{0}$, which demonstrates that gradient induced coherences are of the same leading order as the two other contributions to the interband response. As a final remark, regarding now the intraband response, it shall be noted that there is no Berry curvature corrections to the BE (2), at the leading order $f_{n}^{(1)}\sim\xi_{i}^{-1}$. There are however sub-leading corrections to the intraband linear response at order $\xi_{i}^{0}$ which involve the quantum geometry, and are related to the intrinsic side-jump and skew-scattering contributions found with the semiclassical approach [35, 36, 37]. This is beyond the scope of the present letter, since a consistent account requires to go beyond the ladder approximation [59], and will be provided within our general framework in a future work.
Correspondence with the Kubo formula– It is well established that the intraband part of the linear response controlled by the BE (2), in the constant and homogeneous case, is identical to the intraband part of the Kubo formula with the vertex correction in the ladder approximation [15, 60]. In this section, we establish a similar equivalence for the interband response as given by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). According to the Kubo formula at finite temperature in the constant and translationally invariant limits [15, 16], the expectation value of an observable ${\hat{o}}(\bm{p})\equiv{\hat{o}}$ reads $o^{(1)}=-eE^{i}\int_{p}\lim_{\omega\rightarrow 0}R^{i}(\omega)/({\bf i}\omega)$, with $\int_{p}\equiv\int\frac{{\rm d}^{d}{{p}}}{(2\pi\hbar)^{d}}$ and with $R^{i}({\bm{p}},\omega)\equiv R^{i}(\omega)$ being the analytical continuation (AC) of the Matsubara response function i.e., $R^{i}(\omega)={\cal R}^{i}(\omega_{\nu}\rightarrow-{\bf i}(\hbar\omega+{\bf i}0^{+}))$, with
${\cal R}^{i}(\omega_{\nu})=\sum_{n,m,\varepsilon_{\mu}}{\rm Tr}\left[\hat{P}_{n}{\hat{o}}\hat{P}_{m}\hat{\mathcal{G}}(\varepsilon_{\mu}+\omega_{\nu}){\hat{J}}^{i}\hat{\mathcal{G}}(\varepsilon_{\mu})\right],$ (5)
with $\sum_{\epsilon_{\mu}}\equiv(1/\beta)\sum_{\varepsilon_{\mu}}$ and $\hat{\mathcal{G}}({\bm{p}},\varepsilon_{\mu})\equiv\hat{\mathcal{G}}(\varepsilon_{\mu})$ the matrix-valued Matsubara Green’s function, while $\varepsilon_{\mu}$ and $\omega_{\nu}$ are fermionic and bosonic frequencies and ${\hat{J}}^{i}(\bm{p})\equiv{\hat{J}}^{i}$ is the disorder-dressed current vertex. The Bethe-Salpeter equation for this vertex, in the ladder approximation, is
${\hat{J}}^{i}={\hat{j}}^{i}+n_{i}v_{0}^{2}\int_{\underaccent{\tilde}{p}}\underaccent{\tilde}{\hat{\mathcal{G}}}(\epsilon_{\mu}+\omega_{\nu})\underaccent{\tilde}{\hat J}^{i}\underaccent{\tilde}{\hat{\mathcal{G}}}(\epsilon_{\mu}),$ (6)
with $^{\hat{j}}^{i}=\partial_{p^{i}}\hat{h}$ the bare current vertex. From the corresponding expression of $o^{(1)}$ in terms of the density matrix, we immediately establish the correspondance
$\!\!{\hat{\rho}}^{(1)}_{nm}=-eE^{i}\!\lim_{\omega\rightarrow 0}\frac{\left[\sum_{\epsilon_{\mu}}\!\!\hat{P}_{n}\hat{\mathcal{G}}(\epsilon_{\mu}+\omega_{\nu}){\hat{J}}^{i}\hat{\mathcal{G}}(\epsilon_{\mu})\hat{P}_{m}\right]_{\scriptscriptstyle AC}}{{\bf i}\omega},$ (7)
in which the AC is being performed before taking the $\omega\rightarrow 0$ limit. The Matsubara Green’s function $\hat{\mathcal{G}}(\epsilon_{\mu})$, in presence of disorder, can be obtained from the retarded (advanced) Green’s functions $\hat{g}^{R(A)}({\bm{p}},\varepsilon)\equiv\hat{g}^{R(A)}$, which corresponds to its analytical continuation in the upper (lower) complex plane. Introducing the decomposition $\hat{g}^{\scriptscriptstyle R(A)}=\sum_{n}g^{\scriptscriptstyle R(A)}_{n}\hat{P}_{n}+\sum_{n\neq m}\hat{P}_{n}\hat{g}^{\scriptscriptstyle R(A)}_{nm}\hat{P}_{m}$, and at the level of the self-consistent Born approximation, we have $g^{\scriptscriptstyle R(A)}_{n}=[\varepsilon-\varepsilon_{n}\begin{subarray}{c}\scriptscriptstyle+\\ \scriptscriptstyle(-)\end{subarray}{\bf i}\hbar/(2\tau_{n})]^{-1}$, with $\frac{1}{\tau_{n}}=\int_{\underaccent{\tilde}{p}}\sum_{m}\delta(\varepsilon_{n}-\underaccent{\tilde}{\varepsilon}_{m})W_{nm}$, while
$\!\!\!\hat{g}^{\scriptscriptstyle R(A)}_{nm}=\begin{subarray}{c}\scriptscriptstyle+\\ \scriptscriptstyle(-)\end{subarray}2{\bf i}\pi n_{i}v_{0}^{2}g^{\scriptscriptstyle R(A)}_{n}g^{\scriptscriptstyle R(A)}_{m}\sum_{q}\!\!\int_{\underaccent{\tilde}{p}}\!\delta(\varepsilon-\underaccent{\tilde}{\varepsilon}_{m})\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\hat{P}_{m}.$ (8)
We can also write ${\hat{J}}^{i}=\sum_{n}J^{i}_{n}\hat{P}_{n}+\sum_{n\neq m}\hat{P}_{n}{\hat{J}}^{i}_{nm}\hat{P}_{m}$. At leading order in $\xi_{i}$, we then obtain the band projected part of Eq.(6) as
$J_{n}^{i}=j_{n}^{i}+\theta_{\mu\nu}\int_{\underaccent{\tilde}{p}}\sum_{m}\delta(\varepsilon-\underaccent{\tilde}{\varepsilon}_{m})\underaccent{\tilde}{\tau}_{m}\underaccent{\tilde}{J}^{i}_{m}W_{nm},$ (9)
and its interband components as
$\!\!\!{\hat{J}}^{i}_{nm}\!=\!{\hat{j}}^{i}_{nm}+\frac{2\pi n_{i}v_{0}^{2}}{\hbar}\theta_{\mu\nu}\int_{\underaccent{\tilde}{p}}\sum_{q}\delta(\varepsilon-\underaccent{\tilde}{\varepsilon}_{q})\underaccent{\tilde}{\tau}_{q}\underaccent{\tilde}{J}_{q}^{i}\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\hat{P}_{m},$ (10)
with $\theta_{\mu\nu}\equiv\theta(-\epsilon_{\mu}(\epsilon_{\mu}+\omega_{\nu}))$, by using the analytic structure of the Matsubara Green’s function [43]. From Eq.(7) with $n=m$, one obtains at the same leading order $f_{n}^{(1)}=-\partial_{\varepsilon}f_{n}^{(0)}\tau_{n}\sum_{i}(eE^{i})J^{i}_{n}$, which confirms the expected equivalence between Eq.(9) and Eq.(2), and provides an explicit expression for $J^{i}_{n}$ in terms of $f_{n}^{(1)}$. Then, going back to Eq.(7), it is clear that the interband components of the density matrix are either contributed by the interband components of the current vertex or by those of the two Green’s functions. Then, the first term on the RHS of Eq.(10) gives the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). As for the third term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), the second term of Eq.(10)) yields the terms proportional to $\underaccent{\tilde}{f}^{(1)}_{q}$, whereas the interband components of the Green’s functions, as derived from Eq.(8), lead to the terms proportional to $f^{(1)}_{n}$ and $f^{(1)}_{m}$ [43]. This completes the demonstration of the equivalence between our formalism and the Kubo formula in the ladder approximation, in the constant and homogeneous limits. One needs of course to add that our formalism also validly captures the part of the linear responses slowly varying in time and space at the same leading orders in the semiclassical and disorder expansions, in the case of finite mesoscopic systems. This is much more challenging to achieve with fully quantum calculations.
Pedagogical application to SOT in the 2DEG– As a pedagogical example, we address now the microscopic theory of SOT in the weakly disordered 2DEG, a topic previously investigated with the Kubo formula in the ladder approximation in [20]. The model Hamiltonian is
$\hat{h}({\bm{p}})=\varepsilon_{0}({\bm{p}})\hat{\sigma}_{0}+J_{e}{\bm{m}}\cdot\hat{\bm{\sigma}}+{\bm{b}}_{\scriptscriptstyle SO}({\bm{p}})\cdot\hat{\bm{\sigma}},$ (11)
with ${\bm{p}}\equiv(p^{x},p^{y})$, $\varepsilon_{0}$ some paramagnetic band dispersion, $J_{e}>0$ the exchange constant, ${\bm{m}}$ a classical unit vector aligned with the magnetization and ${\bm{b}}_{\scriptscriptstyle SO}$ some momentum dependent SO field with $b_{\scriptscriptstyle SO}\ll J_{e}$. The operator $\hat{\bm{\sigma}}$ is the vector of Pauli matrices, while $\hat{\sigma}_{0}$ is the identity matrix in spin space. The Hamiltonian (11) is a generalization, to arbitrary band and SO field dispersions, of the model considered in [20]. The dynamics of ${\bm{m}}$ is controlled by the Landau-Lifschitz-Gilbert equation [61], which in the dissipationless limit and spatially uniform case reads $(M_{\scriptscriptstyle S}/\gamma)\partial_{t}{\bm{m}}={\bm{m}}\times(\delta\mathcal{F}/\delta{\bm{m}})\equiv{\bm{T}}$, with $\gamma(<0)$ the gyromagnetic ratio, $M_{\scriptscriptstyle S}$ the saturation magnetization, $\mathcal{F}$ the thermodynamic potential and ${\bm{T}}$ the torque density. The SOT, in linear response to a quasi uniform and constant electric field, is then given by ${\bm{T}}^{(1)}={\bm{m}}\times(\delta\mathcal{F}^{(1)}/\delta{\bm{m}})$, with $\mathcal{F}^{(1)}$ the linear change in $\mathcal{F}$. By application of the Hellmann-Feynman theorem [62, 63] at finite temperature [64], we then derive ${\bm{T}}^{(1)}={\bm{m}}\times\langle\delta\hat{h}/\delta{\bm{m}}\rangle^{(1)}=J_{e}{\bm{m}}\times\langle\hat{\bm{\sigma}}\rangle^{(1)}$. The band energies and associated projectors for the Hamiltonian (11) are trivially obtained as $\varepsilon_{\sigma}=\varepsilon^{0}+\sigma\Delta$ and $\hat{P}_{\sigma}=\frac{1}{2}\left(\hat{\sigma}_{0}+\sigma{\bm{u}}\cdot\hat{\bm{\sigma}}\right)$, respectively ; in which we have introduced $\sigma\in\{+1,-1\}\equiv\{\uparrow,\downarrow\}$, with $\Delta\approx J_{e}(1+{\bm{m}}\cdot\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO})$ and ${\bm{u}}\approx{\bm{m}}+{\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$, as we are approximating from now on all quantities at leading order in $\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}={\bm{b}}_{\scriptscriptstyle SO}/J_{e}$. The interband Berry connection reads $\hat{\mathcal{A}}^{i}_{\sigma\underaccent{\tilde}{\sigma}}\ \approx\ \frac{{\bf i}\underaccent{\tilde}{\sigma}}{4}\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp}\cdot\hat{\bm{\sigma}}\ +\ \frac{1}{4}({\bm{m}}\times\partial_{p^{i}}\underaccent{\bar}{\bm b}_{\perp})\cdot\hat{\bm{\sigma}}$, with $\underaccent{\bar}{\bm b}_{\perp}={\bm{m}}\times(\underaccent{\bar}{\bm b}_{\scriptscriptstyle SO}\times{\bm{m}})$. The intrinsic contribution to the SOT, from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), immediately ensues as
${\bm{T}}^{(1)}_{int}\approx\frac{\hbar}{2}\int_{p}\left(f^{(0)}_{\uparrow}-f^{(0)}_{\downarrow}\right)(e{\bm{E}})\cdot\partial_{\bm{p}}\underaccent{\bar}{\bm b}_{\perp}.$ (12)
After recognizing that ${\bm{E}}\cdot\partial_{\bm{p}}\underaccent{\bar}{b}_{\perp}^{i}=\partial_{\bm{p}}\cdot(\underaccent{\bar}{b}_{\perp}^{i}{\bm{E}})$ and assuming the system to be degenerate, an application of the divergence theorem to the RHS of Eq.(12) yields
${\bm{T}}^{(1)}_{int}\approx\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ \frac{(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{0}}{|\partial_{\bm{p}}\varepsilon_{0}|}\underaccent{\bar}{\bm b}_{\perp},$ (13)
with $dl_{p}$ being the infinitesimal line element along the 1D Fermi surface for band $\sigma$ in absence of SO interaction i.e., $FS_{\sigma}$ 111 In the case of a Fermi surface intersecting the Brillouin zone (BZ) boundary, the line integrals over some portions of this boundary resulting from the application of the divergence theorem does not contribute to the RHS of Eq.(12). Indeed, since the 2D BZ is by definition a primitive unit cell of the reciprocal space over which $\hat{h}$ is periodic, such contributions will cancel out by pairs, from pairs of line segments on opposite sides of the BZ boundary, one being the translation of the other by a primitive reciprocal lattice vector. Hence, we have established that under the stated hypothesis, the intrinsic SOT, which according to Eq.(12) would be classified as a Fermi sea contribution, can equally be expressed as a Fermi surface one according to Eq.(13). This exactly parallels the classical result obtained by Haldane in the case of the non quantized part of the AHE conductivity [44]. Having established this non trivial result, we focus now on the disorder correction to ${\bm{T}}^{(1)}_{int}$, which is stemming from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Specializing it to the case at hand, we readily obtain this contribution to the interband Wigner function as
$\!\!\!\!\hat{\rho}^{(1)}_{dis}\!\approx\!\frac{\pi n_{i}v_{0}^{2}}{2J_{e}}{\bm{m}}\times\!\sum_{\sigma,\underaccent{\tilde}{\sigma}}\underaccent{\tilde}{\sigma}\!\!\!\!\int\displaylimits_{{FS}_{\scriptscriptstyle\underaccent{\tilde}{\sigma}}}\!\!\!\!\!\frac{dl_{\underaccent{\tilde}{p}}}{(2\pi\hbar)^{2}}\!\frac{f_{\sigma}^{\scriptscriptstyle(1)}\!-\!\underaccent{\tilde}{f}_{\underaccent{\tilde}{\sigma}}^{\scriptscriptstyle(1)}}{|\partial_{\underaccent{\tilde}{\bm p}}\underaccent{\tilde}{\varepsilon}_{0}|}(\underaccent{\bar}{\bm b}_{\perp}\!-\underaccent{\tilde}{\underaccent{\bar}{\bm b}}_{\perp})\cdot\hat{\bm{\sigma}},$ (14)
with the distribution function to be derived without SO interaction. In this limit, the scattering Kernel $W_{\sigma,\underaccent{\tilde}{\sigma}}$ becomes diagonal and the BEs for the two spin bands decouple. At this point, we further specialize the problem to a parabolic 2DEG as in [20] i.e., to $\varepsilon_{0}\equiv p^{2}/(2m^{\star})$ with $m^{\star}$ the effective mass. One can then readily solves the BEs with $f_{\sigma}^{(1)}=\delta\left({\varepsilon}_{\sigma}-\varepsilon_{\scriptscriptstyle F}\right)(\tau_{F}/m^{\star})(e{\bf E})\cdot{\bm{p}}$, with the relaxation time and the density of states at the Fermi level respectively given by $1/\tau_{F}=(2\pi n_{i}v_{0}^{2}D_{F})/\hbar$ and $D_{F}=m^{\star}/(2\pi\hbar^{2})$. Substituting this into Eq.(14), and computing the associated spin density, leads to
${\bm{T}}^{(1)}_{dis}\approx-\frac{\hbar}{2}\sum_{\sigma}\sigma\int\displaylimits_{FS_{\sigma}}\!\!\frac{dl_{p}}{(2\pi\hbar)^{2}}\ (e{\bm{E}})\cdot\frac{\bm{p}}{p}\ \underaccent{\bar}{\bm b}_{\perp},$ (15)
exactly canceling the intrinsic torque as given by Eq.(13). This generalizes, to arbitrary weak SO interaction, the central result obtained in [20] with considerable more efforts while relying on a diagrammatic expansion of the Kubo formula in the ladder approximation, for an SO interaction restricted to Rashba or Dresselhaus types.
Conclusions – We have provided a quantum kinetic formulation of the linear response to a an electric field slowly varying in time and space, for a general multiband electron system in presence of weak scalar disorder. While our formalism constitutes a theoretical and computational alternative to the Kubo formula in the stated context, with distinct advantages for finite systems and in capturing disorder effects beyond the CBA, we have proven the equivalence of the two formulations at the ladder approximation level in the homogeneous case. As a pedagogical example, we applied our approach to SOT theory for a simple two-band model of 2DEG amenable to analytical calculations, recovering and generalizing with striking efficiency a major vertex correction, while uncovering a new instance of intrinsic response as a Fermi surface property. We shall stress that our formalism can be applied with equal effectiveness to more involved multiband models up to density functional theory generated tight-binding Hamiltonian, however requiring in such case a numerical resolution of the Boltzmann equation controlling the intraband response. We are already undertaking such studies, for instance to shed new light on the influence of disorder on the SOT in van der Waals ferromagnets exhibiting anisotropic band structure with $C_{3}v$ symmetry [66], a subject of high current interest [18]. It shall also be noted that the disorder model can be easily extended through the consideration of proper self-energies, for instance to include SO scattering contributions. Finally, we want to mention some potential important implications of our formalism on another timely topic i.e., the theory of out-of-equilibrium orbital angular momentum (OAM) in time reversal and inversion symmetric metals, as a possible common explanation for several recently observed effects [47, 45, 46]. In this context, it is to be noted that most Kubo formula based studies have focused so far on the computation of the intrinsic OHE conductivity in the bulk [14, & references therein], irrespective of the chosen definition for the OAM observable. In our formulation, the intrinsic OHE current can be recovered from the first term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). Beyond the possible fragility of this current under disorder corrections [21], which can be straightforwardly assessed from the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we want to point out the importance of the second term i.e., of quantum coherences generated by the gradients in the distribution functions. It can be shown [67] that this is a source of orbital edge accumulation, that is only partially correlated to the OHE conductivity through the Berry connection and which may be an important explanatory factor to the previously cited experimental observations.
Acknowledgements.
We thank Mairbek Chshiev, Henri Jaffrès, Sergey Nikolaev, Frédéric Piéchon and Xavier Waintal for fruitful discussions. One of the author (TV) acknowledges the financial support of the Laboratoire Albert Fert (UMR137) CNRS, Thales, Université Paris-Saclay, France.